# Towards Architecture-Agnostic Untrained Network Priors for Image Reconstruction with Frequency Regularization

### v2

@@ 3.2 Architecture-agnostic Frequency Regularization
where $k_{\ell}$ is a learnable Lipschitz constant for each layer, $\|\cdot\|_{\text{p}}$ is chosen as the $\ell_{\infty}$ norm and $\mathtt{SoftPlus}(k_{\ell})=\text{ln}(1+\text{exp}(k_{\ell}))$ ensures the learned Lipschitz bounds are non-negative. Such formulation only normalizes $W_{\ell}$ if its matrix norm is larger than the learned Lipschitz constraint during training. Integrating the ultimate Lipschitz regularization with learned Lipschitz constants into Eq. 2, our regularized training objective is:
@@ 4.1 Setup and datasets
We first (1) validate the proposed methods’ effectiveness in enhancing the performance of untrained networks across various architectural configurations, especially those originally underperforming. We then (2) benchmark the enhanced versions of those compact architectures against established supervised and self-supervised methods on both in-domain and out-of-domain datasets in terms of accuracy and efficiency. We also (3) compare our methods with self-validation-based early stopping [48] on overcoming overfitting, and show that it is complementary to our approach by further shortening the reconstruction time. Finally, we demonstrate the utility of our methods in (4) general image inpainting and denoising tasks and perform spectral bias analysis on all evaluated tasks.
The MRI experiments were performed on two publicly available datasets including the multi-coil knee and brain MRI images from fastMRI database [25], and multi-coil knee MRI images from Stanford 3D FSE knee dataset [19]. The fully-sampled $k$-space data was retrospectively masked by selecting 25 central $k$-space lines along with a uniform undersampling at outer $k$-space, achieving the standard 4×4$\times$4 × acceleration. For training the supervised baseline, the knee training set consists of 367 PD and PDFS slices and the brain training set consists of 651 slices with a mixture of T1, T2 weighted images. 50 knee slices and 50 brain slices were sampled from the respective multi-coil validation datasets for evaluation.
@@ 4.2 Implementation Details
Without loss of generality, the base architectures considered in our work are N-level encoder-decoder architectures with full skip connections. The architectures are isotropic with the same width and kernel size throughout the layers. All evaluated architectures are trained for 3k iterations using mean absolute error and Adam optimizer [22] with a learning rate of 0.008. Unless otherwise specified, the results at the last iteration are reported. The input is drawn from a uniform distribution $z\sim\mathcal{U}(0,1)$. The filter size of the Gaussian blur was set to 5 and the sigma value was randomly sampled from $[0.5,2.0]$ for every slice. $M$ and $\beta$ for the Kaiser-based upsamplers are chosen to be $\{15\times N-1,5\}$ and $\{5\times N\}$ for knee data ($N$ denotes the nth-level), and $\{5\times N\}$ and $\{5\times N\}$ for brain data, respectively. $\lambda$ is set to $1$ for Lipschitz regularizer.
@@ 4.3 Effectiveness in reducing architectural sensitivity
Fig. 5 gives a quantitative overview of the substantial improvement using our approach in architectures with diverse configurations on both knee and brain datasets. The different results of the original architectures also confirm the influences of architectural choices on performance. Notably, before applying our methods, the deeper and narrower architectures tend to perform better than their counterparts (more in appendix). This trend aligns with previous works [9, 3, 10, 45] where these architectures tend to be favored in inpainting-like tasks. Here we identify particularly their counterparts (i.e., $\mathbf{A_{x}\mathunderscore_{256}}$ ) as "underperforming" architectures. As will be shown in our spectral bias analysis in Sec. 4.5 and appendix, these underperforming architectures learn high frequencies more quickly (though this may be desired for other tasks [30]) and are more susceptible to overfitting, incurring severe artifacts in the output (Figs. 6 and 9). When applied with our methods, as detailed in Tab. 2 and Tab. 3, a large boost in performance is observed in all architectures, especially in $\mathbf{A_{2}\mathunderscore_{256}}$.
We observe that using low-passed inputs via either selected Fourier features or blurring brings the most benefits to the shallower architectures. Better results are achieved when combined with Lipschitz regularization on the layers. On the other hand, deeper architectures benefit more from the Kaiser-based upsampler, which can be seen as performing low-pass filtering on the input feature maps within the network, beyond the initial input layer. We further note that the hyperparameters required for upsampling differ between knee and brain data (Sec. 4.2), with the knee data requiring greater attenuation. This is also consistent with previous findings that they require different numbers of channels for the architecture [9]. Our methods greatly alleviate the need for such architectural tuning by instead allowing for the adjustment of a few key hyperparameters.
@@
Table 5: Out-of-domain evaluation among supervised and untrained methods and comparisons with self-validation-based early stopping[48].
Method In-domain Out-domain Runtime (mean±$\pm$±std) PSNR SSIM PSNR SSIM Train Inference Trained U-Net 31.16 0.776 29.16 0.724 $\sim$1.5 days $0.1\pm 0.003$ sec Untrained CS-$\ell_{1}$ [18] 28.23 0.633 22.46 0.407 ZS-SSL [48] 32.00 0.773 31.74 0.805 $26.1\pm 3.5$ mins DIP [45] 29.16 0.628 28.89 0.664 $9.2\pm 0.3$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ 27.62 0.575 26.03 0.550 $5.5\pm 0.1$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ w. Self-Val 29.59 0.695 27.59 0.641 $0.2\pm 0.2$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ (Ours) 32.07 0.781 31.43 0.790 $6.4\pm 0.4$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ (Ours) w. Self-Val. 31.97 0.776 31.30 0.800 $0.3\pm 0.1$ mins
@@ 4.4 Benchmark results
We then compare our regularized network priors with several established MRI reconstruction methods, including a supervised baseline, a state-of-the-art self-supervised method (ZS-SSL) [48] and other untrained networks (ConvDecoder [9], Deep Decoder [14]). We adopt two lightweight architectures with distinct configurations as the base model. Visual comparisons are shown in Fig. 7.
Our methods enable the previously underperforming architectures to match the performance of ZS-SSL and the supervised UNet on in-domain fastMRI knee data (Tab. 4) and surpass the trained UNet on out-of-domain Stanford 3D FSE data (Tab. 5) as untrained methods are agnostic to domain shifts. The datasets for training the UNet contain 367 knee slices and 651 brain slices, respectively, covering various kinds of contrast weightings. Our enhanced networks also clearly outperform other untrained networks and the compressed-sensing (CS) based reconstruction. These results demonstrate the benefits of enhancing the untrained methods for generalizable reconstruction. ZS-SSL is an unrolling-based method where the network, i.e., a ResNet [44], is adopted as a denoiser. Comparisons with its UNet-variant, i.e.,ZS-SSL-UNet, are included in the appendix. Additionally, ZS-SSL uses a self-validation strategy for early stopping. We further compare it with our methods. Tab. 5 shows that self-validation is not as effective as ours. This is because, while self-validation mitigates overfitting by stopping near the peak PSNR, it does not fundamentally address the issues of underperforming architectures, i.e., the peak PSNR remains subpar. Our methods alleviate overfitting while generally enhancing the peak performance (e.g., Fig. 3), which we further analyze in the next section. Hence, self-validation-based early stopping is complementary to our goal of reducing the architectural sensitivity, and can be integrated into our approach to further shorten the reconstruction time as shown in Tab. 5. Overall, our enhanced network priors achieve comparable performance than ZS-SSL while being orders of magnitude faster.
@@ 4.5 Spectral bias analysis
Figure 8: Measurement of the spectral bias. Our methods allow high frequencies to be fitted at a later iteration, alleviating overfitting and generally beneficial to inpainting-like tasks (i.e., MRI) with higher peak performance. PSNR shown for $\mathbf{A_{2}\mathunderscore_{256}}$.
To examine the efficacy of the proposed methods on frequency control over the network, we measure the spectral bias using the metric— frequency-band correspondence (FBC) [41] that calculates the element-wise $|\mathcal{F}(x)|/|\mathcal{F}(y)|$ between the output $x$ and target image $y$, categorizes it into five frequency bands radially and computes the per-band averages. Higher values indicate higher correspondence. We trace the evolutions of FBC for $\mathbf{A_{8}\mathunderscore_{256}}$ and $\mathbf{A_{2}\mathunderscore_{256}}$ and the corresponding PSNR curves throughout the training iterations in three different tasks.
As shown in Fig. 8, the original architectures tend to readily fit all frequencies, especially $\mathbf{A_{2}\mathunderscore_{256}}$, corresponding to its worst performance among all compared architectures. Our methods delay the convergence of higher frequencies in both architectures across all three tasks as designed, which generally leads to higher peak performance for inpainting-like tasks. We hypothesize that a bias towards lower frequencies helps the model leverage spatial information for improved inter/extrapolation capabilities. Qualitative results in Fig. 9 validate our hypothesis. Given that MRI reconstruction resembles an inpainting task for $k$-space measurements, we expect similar improvement using our regularized network priors, as proved in our experiments. Note that although a stronger spectral bias may not always be optimal, the hyperparameters of our methods enable flexible adaptations for the specific task (more experiments in the appendix).
Figure 9: Inpainting and denoising ($\sigma=25$) experiments for demonstrating the improvement on underperforming architectures.
@@ 5 Conclusion
We introduce efficient, architecture-agnostic methods for frequency control over the network priors, offering a novel solution to simultaneously address the key challenges present in untrained image reconstruction. Our approach requires only minimal modifications to the original DIP scheme while achieving significant gains in accruacy and efficiency as evidenced in MRI reconstruction and natural image restoration tasks, making it a stronger zero-shot baseline reconstructor with the potential for seamless integration with other advancements in self-supervised reconstruction.
Appendix
A Natural image experiments
A.1 Inpainting
To the best of our knowledge, our method is the first attempt that addresses challenges related to architecture, overfitting and runtime simultaneously.
To compare with prior DIP methods, we 1) first employ the network configurations that work best in their respective settings [7, 41], i.e., the architectures used are different in different methods, as shown in Fig. 10 and Fig. 11, we then 2) use an underperforming architecture for all competing methods, as shown in Fig. 12.
Figure 10: Qualitative comparisons with previous DIP methods on inpainting.
Figure 11: Qualitative comparisons with previous DIP methods.
Figure 12: When unsuitable architectures are used, e.g., UNet w. deconvolutions, ResNet w/o. upsampling, SGLD [7] fails to perform restoration. This confirms the importance of architectural decisions in DIP, and that previous methods do not address the architecture-related challenges. Note that deconvolutions have been reported to be not suitable for DIPs [30, 15]. Similarly for ResNet, which does not have any upsampling layers [30].
A.2 Denoising
We first show the results when each method operates in its original setting (Fig. 13), and then evaluate them when an unsuitable architecture is used (Fig. 14).
Figure 13: Qualitative comparisons with previous DIP methods on denoising. Ours trades off the metrics for sharpness.
Figure 14: Denoising results. The base network used in each competing method is replaced by an underperforming architecture, i.e., $\mathbf{A_{2}\mathunderscore_{256}}$.
A.2.1 Transformer.
Figure 15: Qualitative denoising results of a transformer [4]. Our method substantially alleviates the overfitting and enhances the peak PSNR.
Besides CNN, we show here the result on Swin U-Net [4], which consists of only Swin Transformer blocks and skip connections, i.e., no upsampling is involved. As noted in a recent study [30], the unlearnt upsampling is the driving force behind the spectral bias of DIP, and such transformers are more difficult to perform denoising. This finding assumes the white noise as the input. Here, we show that constraining the bandwidth of the noise input enables long-lasting denoising even in such models.
B Influences of hyperparameters
As our methods contain several hyperparameters, we visualize their influences on the frequency control and hence the regularization effects over the output image in Fig. 16.
$s$ and $\sigma$ are associated with the Gaussian blur kernel applied on the noise input. The larger the $s$ and/or $\sigma$ is, the more high frequencies are removed from the noise input (i.e., the smoother the output is). $M$ and $\beta$ are for adjusting the attenuation extents of the Kaiser-based upsamplers. The higher the $M$ and/or $\beta$ is, the larger the attenuation of the high-frequency replica (i.e., the smoother the output is).
Figure 16: Visualizations of the frequency control with our methods in denoising experiments.
C Comparisons with early-stopping
Figure 17: Comparisons with self-validation-based early stopping. Although early stopping alleviates overfitting, it cannot fundamentally improve the underperforming architectures. Our methods can improve their peak performance while also mitigating overfitting.
To complement Table 5 in the main text, here we visualize in Fig. 17 that early stopping, even though prevents further performance decay, cannot fundamentally improve the underperforming architectures. In other words, early stopping could not cope with ill-designed architectures.
D Architectural influences
To better inform our method design, we 1) investigate the architectural influences in the context of MRI reconstruction, and also 2) validate our findings on image inpainting and denoising. Our results confirm that the reconstruction outcome is sensitive to basic architectural properties.
D.1 Crucial Architectural Elements
We first pinpointed the core architecture elements that have a critical impact on the performance.
Experimental setup i. Since a decoder is the minimum requirement for reconstruction, we experimented with two types of 7-layered decoder-only architectures, i.e., ConvDecoder [9] and Deep Decoder [14]. Experiments were performed on the 4× under-sampled multi-coil knee MRI from fastMRI database [25].
Figure 18: Influences of architectural elements. Results averaged across three different widths.
Upsampling (interpolation filter). Fig. 18(a) suggests an interesting result: removing the unlearnt upsampling, e.g., bilinear, leads to either failure or unstable results (see gray curves). Unlike transposed convolution, the unlearnt upsampler is essentially a fixed low-pass interpolation filter that attenuates the introduced high-frequency replica and also the signal. Frequency response of bilinear interpolation filter decays more rapidly than that of nearest neighbor as the frequency increases (Fig. 18 (b)), suggesting stronger attenuation and smoothing effects. Hence, bilinear upsampling typically biases the network towards generating smoother outputs, as prevalent in generative models [40]. Transposed convolutions, however, are not guaranteed to be low-passed as they are learnable. Due to the spectral bias of network layers, they may be low-passed during early training to still enable reconstruction, but the results could be unstable (green curves).
Convolutional layers. When the unlearnt upsampling operations are absent, ConvDecoder ($3\times 3$) still enables reconstruction while Deep Decoder ($1\times 1$) fails completely (Fig. 18(a)). A similar phenomenon is also reported in image denoising [5, 30]. This again be attributed to CNN’s inherent spectral bias and suggests that the size of the kernel also matters, further corroborated in Tab. 7.
Discussion. Results of this pilot experiment suggest that the spatial kernels with low-pass characteristics, either learnable or unlearnt, are crucial to the success of untrained network priors. In particular, bilinear upsampling with a fixed low-pass filter produces more stable and better results (blue curves).
D.2 Depth, Width and Skip Connections
Here, we demonstrate that insights gained about the unlearnt upsampling can aid in understanding the connection between architectural characteristics and the reconstruction task.
Experimental setup ii. For this large-scale validation, we experimented with an isotropic encoder-decoder architecture used in the original DIP, i.e., equal width and kernel size for all layers throughout the network. Design choices are detailed in Tab. 6. Experiments were performed on the publicly available 4×4$\times$4 × under-sampled multi-coil knee MRI from fastMRI database [24].
Table 6: Test bed for studying the architectural influences of an encoder-decoder untrained networks.
Archi. Type Depth (d) # of Skips (s) Width (w) Kernel Size (k) $\mathbf{A_{d}\mathunderscore_{s}\mathunderscore_{w}\mathunderscore_{k}}$2-L, 3-L, 4-L, 5-L, 8-L} {zero, half, full} {32, 64, 128, 256} $3\times 3,5\times 5$}
Table 7: Influences of typical architectural design choices in knee reconstruction. Deeper and/or Narrower architectures tend to perform better; skip connections influence the deep architectures more; larger kernels perform slightly better. $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ performs the best (in lime); $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ performs the worst (in red).
Width ($\downarrow$) Depth ($\uparrow$) Archi. PSNR SSIM Archi. PSNR SSIM Archi. PSNR SSIM Archi. PSNR SSIM $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 26.67 0.530 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 27.12 0.543 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 27.70 0.583 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 28.47 0.641 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.22 0.590 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 28.59 0.605 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 28.55 0.616 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 29.25 0.660 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.68 0.617 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 28.95 0.622 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 28.87 0.624 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 29.70 0.671 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.61 0.613 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 28.87 0.615 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 29.33 0.648 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 29.81 0.680 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.98 0.625 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 29.33 0.637 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 29.45 0.651 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 30.04 0.695 Skip Connections (--) Kernel Size ($\uparrow$) $\mathbf{A_{2}\mathunderscore_{half}\mathunderscore_{256}\mathunderscore_{3}}$ 26.91 0.535 $\mathbf{A_{2}\mathunderscore_{zero}\mathunderscore_{256}\mathunderscore_{3}}$ 26.83 0.535 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 26.67 0.530 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{5}}$ 26.98 0.550 $\mathbf{A_{4}\mathunderscore_{half}\mathunderscore_{256}\mathunderscore_{3}}$ 28.55 0.621 $\mathbf{A_{4}\mathunderscore_{zero}\mathunderscore_{256}\mathunderscore_{3}}$ 27.54 0.697 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.61 0.613 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{5}}$ 28.82 0.624 $\mathbf{A_{8}\mathunderscore_{half}\mathunderscore_{256}\mathunderscore_{3}}$ 29.12 0.669 $\mathbf{A_{8}\mathunderscore_{zero}\mathunderscore_{256}\mathunderscore_{3}}$ 28.51 0.609 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.98 0.625 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{5}}$ 29.12 0.634
Figure 19: Generalizability of different architectures on the masked regions.
What do deeper and narrower architectures produce? (Tab. 7). Theoretically, as the number of layers (depth) or channels (width) increases, the ability of the network to learn arbitrarily high frequencies (details, noise) is typically increased [38]. While this is true for width, we have found that the effect on depth turns out to be attenuated by unlearnt upsampling. As evidenced in Fig. 19, deeper architectures typically generate smoother images, exhibiting a stronger preference for low-frequency information, whereas shallower counterparts, even though they have fewer parameters, are more susceptible to noise and overfitting (red arrows). This is more evident when comparing the same architectures with just different upsamplers, where the architectures with bilinear upsampling (stronger attenuation) are less prone to overfitting than the ones using nearest neighbor (NN) upsampling (cyan vs. blue). Hence, it is not merely the number of parameters but the architectural characteristics promoting low frequencies that seem to be the primary reason for the high performance. Note that all these results are only achievable when unlearnt upsampling is involved (gray dashed curves).
Skip connections. Deep architectures with zero skip connection converge more slowly and may lead to over-smoothing as shown in Fig. 19 (red curves). Skip connections greatly alleviate this issue and introduce more details (cyan curves), which we speculate could be due to the "reduced effective up-sampling rate". Yet, excessive skip connections make a deep architecture behave similarly as a shallower one, generating more noise (Fig. 19 right). Overall, they exert a greater influence on deeper architectures ($\mathbf{A_{8}\mathunderscore_{zero}}<\mathbf{A_{8}\mathunderscore_{full}}<\mathbf{A_{8}\mathunderscore_{half}}$) compared to shallower ones ($\mathbf{A_{2}\mathunderscore_{full}}\approx\mathbf{A_{2}\mathunderscore_{zero}}$).
D.3 Validations on image denoising and inpainting.
We reaffirmed our observations above on image denoising and inpainting, as shown in Fig. 20, Fig. 22, Fig. 23 and Fig. 21.
Figure 20: Denoising experiments. (Left) In non-convolutional networks, removing the upsampling hampers the denoising capability, which cannot be compensated by merely adjusting the network to be more under-parameterized. Transposed convolutions result in a more rapid decline in performance than bilinear upsampling. (Right) Convolutional layers alone exhibit certain denoising effects but necessitate early stopping. The showcased image is from the classic dataset Set9 [8].
Figure 21: Deep architectures with zero skip connection converge more slowly, i.e., $\mathbf{A_{8}\mathunderscore_{zero}\mathunderscore_{256}}$
We argue that the understanding of the upsampling and its interactions with other architectural elements can help explain why deeper networks with fewer skip connections converge more slowly, generate smoother outputs and are less prone to overfitting. Concretely, the upsampling operation inserted in-between the decoder layer slows down the generation of high frequencies required for transforming the lower-resolution feature maps into the higher-resolution target image, primarily due to its role as a fixed low-pass filter. As the network depth increases, the degree of smoothness increases (Fig. 22). Skip connections notably accelerate the convergence (Fig. 21) and ameliorate the over-smoothing issue, likely due to the reduced "effective" upsampling rate. All these observations are consistent with our MRI experiments in Sec. 7.
Figure 22: Denoising experiments. Deeper architectures with few or no skip connections tend to generate smoother outputs compared to the shallower ones.
Figure 23: Inpainting experiments. Deeper architectures with few or no skip connections tend to generate smoother predictions for the masked regions than the shallower architectures. Skip connections make deep architectures perform similarly as the shallower ones.
E Comparisons with ZS-SSL-UNet
In the main paper, we have included the results of the ResNet version of ZS-SSL [48]. Here, we constructed a UNet variant of it, dubbed ZS-SSL-UNet. As shown in Tab. 8, the architecture type impacts not only DIP but also deep unrolling networks, and potentially a broader area, which worth future investigations.
Table 8: Quantitative evaluations. Runtime: mean $\pm$ std mins per slice.
[…]

### v3

@@ 3.2 Architecture-Agnostic Frequency Regularization
where $k_{\ell}$ is a learnable Lipschitz constant for each layer, $\|\cdot\|_{\text{p}}$ is chosen as the $\ell_{\infty}$ norm and $\mathtt{SoftPlus}(k_{\ell})=\text{ln}(1+\text{exp}(k_{\ell}))$ ensures the learned Lipschitz bounds are non-negative. Such formulation only normalizes $W_{\ell}$ if its matrix norm is larger than the learned Lipschitz constraint during training. Integrating the ultimate Lipschitz regularization into Eq. 2, our regularized training objective is
@@ 4.1 Setup and Datasets
We first (1) validate the effectiveness of the proposed methods in enhancing the performance of untrained networks across various architectural configurations, especially those originally underperforming. We then (2) benchmark the enhanced versions of those compact architectures against established supervised and self-supervised methods on both in-domain and out-of-domain datasets in terms of accuracy and efficiency. We also (3) compare our methods with self-validation-based early stopping [46] on overcoming overfitting, and show that it is complementary to our approach by further shortening the reconstruction time. Finally, we demonstrate the utility of our methods in (4) general image inpainting and denoising tasks and (5) perform spectral bias analysis on all evaluated tasks.
The MRI experiments were performed on two publicly available datasets: the multi-coil knee and brain MRI images from fastMRI database [23], and multi-coil knee MRI images from Stanford 3D FSE knee dataset [17]. The fully-sampled $k$-space data was retrospectively masked by selecting 25 central $k$-space lines along with a uniform undersampling at outer $k$-space, achieving the standard 4×4$\times$4 × acceleration. For training the supervised baseline, the knee training set consists of 367 PD and PDFS slices and the brain training set consists of 651 slices with a mixture of T1 and T2 weighted images. 50 knee slices and 50 brain slices were sampled from the respective multi-coil validation datasets for evaluation.
@@ 4.2 Implementation Details
Without loss of generality, the base architectures considered in our work are N-level encoder-decoder architectures with full skip connections. The architectures are isotropic with the same width and kernel size throughout the layers. All evaluated architectures are trained for 3k iterations using mean absolute error and Adam optimizer [20] with a learning rate of 0.008. Unless otherwise specified, the results at the last iteration are reported. The input is drawn from a uniform distribution $z\sim\mathcal{U}(0,1)$. The filter size of the Gaussian blur was set to 5 and the sigma value was randomly sampled from $[0.5,2.0]$ for every slice. $M$ and $\beta$ for the Kaiser-based upsamplers are chosen to be $\{15\times N-1,5\}$ and $\{5\times N\}$ for knee data ($N$ denotes the nth-level), and $\{5\times N\}$ and $\{5\times N\}$ for brain data, respectively. $\lambda$ is set to $1$ for the Lipschitz regularizer.
@@ 4.3 Effectiveness in Reducing Architectural Sensitivity
Fig. 5 gives a quantitative overview of the substantial improvement give by our approach in architectures with diverse configurations on both knee and brain datasets. The different results of the original architectures also confirm the influences of architectural choices on performance. Notably, before applying our methods, the deeper and narrower architectures tend to perform better than their counterparts (more in appendix). This trend aligns with previous works [6, 3, 7, 43] where these architectures tend to be favored in inpainting-like tasks. Here we identify particularly their counterparts (i.e., $\mathbf{A_{x}\mathunderscore_{256}}$) as "underperforming" architectures. As will be shown in our spectral bias analysis in Sec. 4.5 and appendix, these underperforming architectures learn high frequencies more quickly (though this may be desired for other tasks [29]) and are more susceptible to overfitting, incurring severe artifacts in the output (Figs. 6 and 10). When applied with our methods, as detailed in Tab. 2 and Tab. 3, a large boost in performance is observed in all architectures, especially $\mathbf{A_{2}\mathunderscore_{256}}$.
We observe that using low-passed inputs via either selected Fourier features or blurring brings the most benefits to the shallower architectures. Better results are achieved when combined with Lipschitz regularization on the layers. On the other hand, deeper architectures benefit more from the Kaiser-based upsampler, which can be seen as performing low-pass filtering on the input feature maps within the network, beyond the initial input layer. We further note that the hyperparameters required for upsampling differ between knee and brain data (Sec. 4.2), with the knee data requiring greater attenuation. This is also consistent with previous findings that they require different numbers of channels for the architectural choices [6]. Our methods greatly alleviate the need for such architectural tuning by instead allowing for the adjustment of a few key hyperparameters.
@@
Table 5: Out-of-domain evaluation among supervised and untrained methods and comparisons with a self-validation-based early stopping strategy [46].
Method In-domain Out-domain Runtime (mean±$\pm$±std) PSNR SSIM PSNR SSIM Train Inference Trained U-Net 31.16 0.776 29.16 0.724 $\sim$1.5 days $0.1\pm 0.003$ sec Untrained CS-$\ell_{1}$ [16] 28.23 0.633 22.46 0.407 ZS-SSL [46] 32.00 0.773 31.74 0.805 $26.1\pm 3.5$ mins DIP [43] 29.16 0.628 28.89 0.664 $9.2\pm 0.3$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ 27.62 0.575 26.03 0.550 $5.5\pm 0.1$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ + Early Stopped 29.59 0.695 27.59 0.641 $0.2\pm 0.2$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ (Ours) 32.07 0.781 31.43 0.790 $6.4\pm 0.4$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ (Ours) + Early Stopped 31.97 0.776 31.30 0.800 $0.3\pm 0.1$ mins
@@ 4.4 Benchmark Results
We adopt two most lightweight architectures as the base models, and compare our regularized network priors with several established MRI reconstruction methods, including a supervised baseline, a state-of-the-art self-supervised method based on unrolling networks (ZS-SSL) [46] and underparameterized untrained networks (ConvDecoder [6], Deep Decoder [12]). Visual comparisons are presented in Fig. 7.
Comparisons with state-of-the-arts. Our method enables the previously underperforming architectures to match the performance of ZS-SSL and that of the supervised UNet on fastMRI knee data (Tab. 4), and surpass the trained UNet on out-of-domain Stanford 3D FSE data (Tab. 5), demonstrating their advantages in generalizable reconstruction. Our enhanced networks also clearly outperform other lightweight untrained networks, i.e., ConvDecoder and Deep Decoder, which are designed to prevent overfitting. ZS-SSL is an unrolling method where the network, i.e., a ResNet [42], is adopted as a denoiser. Our enhanced network priors achieve comparable performance than ZS-SSL while being orders of magnitude faster, thanks to the more efficient base models.
Comparisons with early stopping (ES). ZS-SSL uses a self-validation strategy for early stopping whereas our method does not necessitate ES.
Figure 8: 8×8$\times$8 × undersampling
Our method not only alleviates overfitting but more importantly, enhances the inter/extrapolation capabilities of the underperforming architectures, leading to higher peak performance than the original scheme (Tab. 5). This cannot be achieved even with the best ES strategy, as ES only halts the training near the peak PSNR for a given architecture, but it cannot fundamentally improve the underperforming architectures whose peak PSNR remains subpar (Fig. 8). We show in Tab. 5 that self-validation based ES can be integrated into our approach to further shorten the reconstruction time from 6 $\rightarrow$0.3 mins. More examples are in Suppl. Fig. 8.
@@ 4.5 Spectral Bias Analysis
Figure 9: Measurement of spectral bias. Underperforming architectures (e.g., $A_{2}\mathunderscore_{256}$, $A_{8}\mathunderscore_{256}$) tend to learn high frequencies hastily, overfit more easily, and extrapolate poorly. Our method effectively mitigates these shortcomings.
To examine how the proposed methods influence the frequency bias of the network, we measure the spectral bias using the metric— frequency-band correspondence (FBC) [39], which first calculates the element-wise $|\mathcal{F}(x)|/|\mathcal{F}(y)|$ between the output $x$ and target image $y$, categorizes it into five frequency bands radially and then computes the per-band averages. Higher values indicate higher correspondence. We trace the evolutions of FBC for $\mathbf{A_{8}\mathunderscore_{256}}$ and $\mathbf{A_{2}\mathunderscore_{256}}$ and the corresponding PSNR curves throughout the training iterations in three tasks.
Fig. 9 shows that the original underperforming architectures tend to fit all frequencies more readily, including high frequencies. This is more evident in $\mathbf{A_{2}\mathunderscore_{256}}$, corresponding to its worst performance among all compared architectures. Our methods substantially delay the convergence of higher frequencies for all three tasks as designed, which decouples the learning of different frequencies. This leads to prolonged denoising effects and enhanced performance for inpainting tasks, including MRI reconstruction, as qualitatively shown in Fig. 10. We speculate that a stronger bias towards lower frequencies helps the model leverage better the spatial information which improves its inter/extrapolation capability. Given that MRI reconstruction resembles an inpainting task for $k$-space measurements, we expect similar improvements in natural image inpainting shown in Fig. 10 with our regularized network priors for $k$-space interpolation in MRI experiments.
Figure 10: Experiments on natural image inpainting and denoising ($\sigma=25$). Our method improves the extrapolation and denoising capabilities of underperforming architectures.
@@ 5 Conclusion
We introduce efficient, architecture-agnostic methods for frequency control over the network priors, offering a novel solution to simultaneously address the key challenges present in untrained image reconstruction. Our approach requires only minimal modifications to the original DIP scheme while achieving significant gains in accruacy and efficiency as evidenced in MRI reconstruction and natural image restoration tasks, making it a stronger zero-shot reconstructior with the potential for seamless integration with other advancements in self-supervised reconstruction.
Acknowledgements
This work was supported in part by the United States National Institutes of Health (NIH) through grants R01CA266702 and R01EB035160.